OpenStudy (anonymous):

Verify the identity. cscx - tan(x/2) = cotx

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ tan\left(\frac{\theta}{2}\right)=\cfrac{1-cos(\theta)}{sin(\theta)}}}\\ \quad \\ \quad \\ csc(x)-tan\left(\frac{x}{2}\right)\implies \cfrac{1}{sin(x)}-\cfrac{1-cos(x)}{sin(x)}\)

OpenStudy (zzr0ck3r):

I didnt see the x/2 and was having a heck of a time.....lol

OpenStudy (jdoe0001):

heh

OpenStudy (anonymous):

thank you